# Opponent Shaping for Antibody Development

### v2

@@ Abstract
Anti-viral therapies are typically designed or evolved towards the current strains of a virus. In learning terms, this corresponds to a myopic best response, i.e., not considering the possible adaptive moves of the opponent. However, therapy-induced selective pressures act on viral antigens to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. To motivate our work, we consider antibody designs that target not only the current viral strains but also the wide range of possible future variants that the virus might evolve into under the evolutionary pressure exerted by said antibodies. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of the viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simple model, we hope that our proposed paradigm will enable the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools.
@@ 1 Introduction
Designing effective therapies to fight off viral pathogens is crucial. However, traditional approaches have been largely myopic, focusing on the current variant of a virus without considering the future evolutionary pressures induced by the therapies themselves. Although this shortsighted approach may yield therapies with high initial efficacy, it fails to account for the virus’s adaptive response, leaving treatments vulnerable to viral escape and the emergence of new therapy-resistant strains [1, 2, 3, 4, 5].
@@
The COVID-19 pandemic starkly illustrated this challenge. While the rapid development of vaccines was a remarkable achievement, concerns quickly arose about their long-term efficacy against new emerging COVID variants [6, 7]. For example, the B.1.351 variant demonstrated that the vaccine loses its effectiveness against new strains [8]. This underscores the need for approaches that consider both the current and future efficacy of a designed therapy.
Toward this goal, we propose to design therapies that target not only current viral variants but also anticipate and target future ones using a model of evolutionary adaptation. Specifically, viruses evolve in response to the selective pressures imposed by our therapies implying that the therapies we design exert a degree of influence over viral evolution. Recognizing this, our goal is to develop therapies which steer the viral evolution away from escape.
@@
This concept aligns closely with the principles of opponent shaping [9]. Opponent shaping is a multi-agent reinforcement learning framework that allows agents to anticipate and influence the future strategies of other agents in their environment. This approach, exemplified by methods such as Learning with Opponent-Learning Awareness (LOLA) [9] and Model-Free Opponent Shaping (M-FOS) [10], encourages agents to consider not only their immediate rewards but also the consequences of their actions on their opponents’ future behaviour. In the context of viral therapy, we view the virus as an adversary that adapts to our chosen therapy strategy. By applying opponent shaping techniques to this problem, we suggest optimising therapeutics to be more robust to future viral mutations and shape viral evolution in favourable directions.
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game which we call the virus-antibody game. Importantly, we include key functional constraints of the antibody and the virus in the game. Our framework applies opponent shaping techniques to optimise antibodies for long-term effectiveness against evolving viral populations, limiting viral escape.
Viral escape occurs when mutations in the virus allow it to successfully evade the host’s immune system [11]. In the context of our game, the virus reduces or eliminates the antibody’s ability to bind effectively, i.e. it finds an effective best-response strategy to the current antibody. We construct a simulated model of viral escape that captures the adaptive responses to antibody pressure, using the Absolut! framework [12] to approximate the binding strength of protein-protein interactions. Importantly, our simulations reveal that the conventionally “optimal” choice of antibody — one that myopically maximises binding to the current virus — is suboptimal when considering a longer time horizon of viral antigen mutations.
Based on this insight, our algorithm instead optimises antibodies with respect to their expected fitness over the entire viral escape curve. We term these optimised antibodies shapers, in contrast to myopic antibodies that are selected based solely on their efficacy against the current viral variant. Our results demonstrate that shaper antibodies not only outperform myopic antibodies in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding.
Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method.
Finally, we present an analysis of the key features that distinguish shaper antibodies from their myopic counterparts. This explainability component provides insights into the amino acid sequence characteristics that contribute to long-term antibody efficacy and viral evolution shaping within our simulations.
In summary, this work introduces a novel and general antibody design framework that leverages opponent-shaping principles. By anticipating and influencing expected viral evolution, our method improves the long-term efficacy of antibodies within our Absolut! simulations, and when combined with more accurate future simulators may assist in the design of antiviral therapies that minimise the rate of viral escape.
@@ 2 Results
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against the current and future virus variants and influence the viral evolution. We approximate the future variants through the Simulated Viral Escape via Evolution (i.e. the inner loop) where the viral antigens are evolved against the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimize, while green arrows represent those they aim to maximize. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [12] and the Miyazawa-Jernigan energy potential matrix [14].
@@ 2.1 Computational Implementation of Opponent Shaping for Antibody Design
To implement opponent shaping for antibody design, we develop a computational framework that models the virus-antibody interaction and optimises antibodies for long-term efficacy. Our approach comprises three main components: defining the virus-antibody game with the player’s payoffs, simulating viral escape, and optimising antibody shapers.
@@ 2.1.1 Virus-Antibody Game
We formalise the interaction between antibodies and viruses as a two-player zero-sum game. In this game, two players – the virus and the antibody – play a game where one player’s gain is the other’s loss. The game is defined by the set of actions available to each player and their respective payoffs.
The players’ actions are represented by their amino acid sequences. The sequences are of an antigen protein for the virus and a fragment of a hypervariable region of the heavy chain for the antibody, specifically the third complementarity-determining region (CDRH3), which is the most important part of the antibody for defining its specificity and affinity [15].
@@
To evaluate the binding strength $B(\cdot,\cdot)$ in our setup, we adapt the Absolut! framework [12]. Absolut! is a simplified simulator for antibody-antigen binding. Unlike sequence-based machine learning models [16, 17, 18, 19] that cannot generalise to novel viral mutations, or computationally intensive molecular dynamics simulations [20], Absolut! offers a fast approach that generalises to both viral and antibody mutations, which is necessary for our application.
@@ 2.1.2 Simulated Viral Escape via Evolution
To simulate the space of possible viral escape trajectories against a chosen antibody, we model the virus learning to play the game through an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate the process of viral escape for a specified horizon of $H$ steps, representing multiple generations of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of the viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods section 4.2 for more details.
We use the antigen protein from the Dengue Virus as the virus for all our experiments, specifically, the structure with PDB code 2R29 [21]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
@@ 2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. We start with a random antibody sequence and generate a population of antibodies of size $P_{a}$ through mutations in the amino acid sequence. Each new antibody in this population is created by introducing a single mutation to the original sequence.
We define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s performance over multiple potential viral escape trajectories of virus $v$:
@@
The optimisation proceeds iteratively for $N$ steps. At each step, we select the antibody with the highest fitness and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape, see Antibody Optimisation Loop in Figure 1a and Methods section 4.3 for more details.
It’s worth noting the computational intensity of this process: a single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for $H$ steps. Consequently, we refer to different optimised antibodies as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length chosen for the viral escape simulations during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. The horizon length significantly impacts the antibody’s long-term effectiveness and the computational resources required for optimisation.
This process can be viewed as meta-optimisation or meta-learning. The inner loop of our optimisation is the viral escape simulation, where we model the virus learning to evade the current antibody. The outer loop or meta-loop is the antibody optimisation process itself, which learns to generate antibodies that perform well across many possible viral escape trajectories. We are designing antibodies that anticipate and influence the escape process of the virus, embodying the core principle of opponent shaping in this biological context.
@@ 2.2 Antibody Shapers Minimise Viral Escape
We validate the effectiveness of the antibody shapers in optimising the escape-averaged antibody fitness function $F^{H}_{v}(a)$ compared to myopic antibodies that only respond to the current virus $v$. The myopic antibodies optimise the fitness function $F^{0}_{v}(a)=R_{a}(v,a)$, which considers only the immediate payoff against the non-mutated virus. For our shaper antibodies, we select a long horizon of $H=100$ to capture extended viral escape trajectories. Both shapers and myopic antibodies are optimised for $N=30$ steps. Figure 2a presents the performance distributions of shapers and myopic antibodies under both objective functions.
@@ 2.2.2 Antibody Shapers with Varying Horizons
However, the number of antibody optimisation steps does not accurately reflect the computational or experimental cost of optimisation. Each simulation of viral escape requires a number of binding samples that increases linearly with the horizon length $H$. Yet, shorter horizon antibodies optimise an objective that diverges further from our antibody objective $F^{100}_{v}(a)$. Consequently, the optimal horizon varies depending on the available computational budget.
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples — queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process — to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths.
Interestingly, $H=20$ shapers perform strongly - nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
We find that the optimisation horizon significantly influences the performance of antibody shapers. While longer horizons generally lead to better long-term performance, the optimal horizon length is dependent on the available computational resources. Thus, it is important to consider the balance between computational cost and the fidelity of the optimisation objective when designing antibodies for long-term effectiveness against evolving viral populations.

### v4

@@ Abstract
Anti-viral therapies are typically designed to target the current strains of a virus. Game theoretically, this corresponds to a short-sighted, or myopic, response. However, therapy-induced selective pressures act on viral antigens to drive the emergence of mutated strains, against which initial therapies have reduced efficacy. Building on a computational model of binding between antibodies and viral antigens (the Absolut! framework), we design and implement a genetic simulation of such viral evolutionary escape. Crucially, this allows our antibody optimisation algorithm to consider and influence the entire escape curve of the virus, i.e. to guide (or “shape”) the viral evolution. This is inspired by opponent shaping which, in general-sum learning, accounts for the adaptation of the co-player rather than playing a myopic best response. Hence we call the optimised antibodies shapers. Within our simulations, we demonstrate that our shapers target both current and simulated future viral variants, outperforming the antibodies chosen in a myopic way. Furthermore, we show that shapers exert specific evolutionary pressure on the virus compared to myopic antibodies. Altogether, shapers modify the evolutionary trajectories of viral strains and minimise the viral escape compared to their myopic counterparts. While this is a simplified model, we hope that our proposed paradigm will enable the discovery of better long-lived vaccines and antibody therapies in the future, enabled by rapid advancements in the capabilities of simulation tools. Our code is available at https://github.com/olakalisz/antibody-shapers.
@@ 1 Introduction
Designing effective therapies to fight off viral pathogens is crucial. However, traditional design approaches only target the current variant of a virus. Although this myopic design approach may yield therapies with high initial efficacy, it fails to account for viral adaptation, leaving treatments vulnerable to becoming ineffective over time [1, 2, 3, 4, 5]. The COVID-19 pandemic starkly illustrated the challenges of adaptive viruses. While the rapid development of vaccines was a remarkable achievement, concerns quickly arose about their long-term efficacy against new emerging COVID variants [6, 7]. For example, the B.1.351 variant demonstrated that the vaccine loses its effectiveness against new strains [8]. This underscores the need for approaches that consider both the current and future efficacy of a designed therapy.
@@
The therapies we design inevitably exert a degree of influence over viral evolution, as the viruses adapt in response to selective pressures imposed by said therapies. Recognizing this, our work aims to develop a method for designing therapies that can influence viral evolution and steer it towards less dangerous variants. To achieve this we utilise principles from opponent shaping [9]. Opponent shaping is a multi-agent reinforcement learning framework that allows agents to anticipate and influence the future strategies of other agents in their environment. This approach, exemplified by methods such as Learning with Opponent-Learning Awareness (LOLA) [9] and Model-Free Opponent Shaping (M-FOS) [10], encourages agents to consider not only their immediate rewards but also the consequences of their actions on their opponents’ future behaviour after a learning update.
Building on these principles, our work focuses specifically on antibody design, framing the interaction between antibodies and surface viral antigens as a two-player zero-sum game. In this game, the antibody’s payoff is primarily determined by its binding strength to the virus, whilst the virus has the opposite payoff. Although our framework can use any binding model (or indeed any sequence-to-function map), in this paper we use our GPU-accelerated implementation of the Absolut! framework [11] to approximate the binding strength of protein-protein interactions. We use this game to model viral escape - the process via which mutations in a virus allow it to successfully evade a host’s immune system [12]. In the language of game theory, this is the virus adapting to the current antibody by repeatedly finding a local best response, decreasing the virus’s binding strength to the antibody. Using this viral escape model we optimise antibodies to be effective over the entire process of viral evolution, resulting in designed antibodies which we refer to as shapers. This is in contrast to only optimising for binding to the initial virus, which results in myopic antibodies.
Importantly, our simulations reveal that myopic antibodies are suboptimal when considering a longer time horizon of viral evolution. Shaper antibodies not only outperform their myopic counterparts in long-term efficacy but also exert a shaping influence on viral evolutionary trajectories, effectively constraining the virus’s ability to escape antibody binding. Our study also explores the trade-offs between the effectiveness of shaper antibodies and the computational resources required for their optimisation, providing insights into the scalability and potential for practical deployment of our method. Finally, we present an analysis of the key features that distinguish between shaper and myopic antibodies.
In summary, this work introduces a novel antibody design philosophy leveraging opponent-shaping principles to improve the long-term efficacy of antibodies against evolving viruses. While our results provide a promising proof of concept, they are based on simplified models of binding and viral escape, limiting direct real-world application. However, as more sophisticated methods for modelling protein interactions and predicting viral evolution emerge, our framework has the potential to significantly impact future antiviral therapy design.
@@ 2 Results
Figure 1: Computational Implementation of Opponent Shaping for Antibody Design. a An overview of our implementation. In the Antibody Optimisation Loop (i.e. the outer loop), we optimise the antibody to perform well against the current and future virus variants and influence the viral evolution. We approximate the future variants through the Simulated Viral Escape via Evolution (i.e. the inner loop) where the viral antigens are evolved against the current antibody over a given horizon length. b The payoffs of the antibody and virus. Red arrows indicate binding interactions that players aim to minimise, while green arrows represent those they aim to maximise. The antibody optimises for binding to the virus while avoiding its anti-target. In this zero-sum game, the antibody’s optimisation indirectly counters the virus’s binding to its target. c Efficient JAX [13] implementation of the binding calculation uses binding poses generated by Absolut! [11] and the Miyazawa-Jernigan energy potential matrix [14].
@@ 2.1 Computational Implementation of Opponent Shaping for Antibody Design
To implement opponent shaping for antibody design, we develop a computational framework that models the virus-antibody interaction and optimises antibodies for long-term efficacy. Our approach comprises three main components: defining the virus-antibody game with the player’s payoffs, simulating viral escape, and optimising antibody shapers. We cover each of these components separately in the following three sections.
@@ 2.1.1 Virus-Antibody Game
We formalise the interaction between antibodies and viruses as a two-player zero-sum game. In this game, two players – the virus and the antibody – play a game where one player’s gain is the other’s loss. The game is defined by the set of actions available to each player and their respective payoffs. The players’ actions are represented by their amino acid sequences. The sequences are of an antigen protein for the virus and a fragment of a hypervariable region of the heavy chain for the antibody, specifically the third complementarity-determining region (CDRH3), which is the most important part of the antibody for defining its specificity and affinity [15].
@@
To evaluate the binding strength $B(\cdot,\cdot)$ in our setup, we adapt the Absolut! framework [11]. Absolut! is a simplified simulator for antibody-antigen binding. Unlike sequence-based machine learning models [16, 17, 18, 19] that cannot generalise to novel viral mutations, or computationally intensive molecular dynamics simulations [20], Absolut! offers a fast approach that generalises to both viral and antibody mutations, which is necessary for our application. However, using Absolut! sacrifices accuracy, a trade-off we discuss more explicitly in Section 3.
@@ 2.1.2 Simulated Viral Escape via Evolution
Viral Escape is the process of a virus adapting to a host’s immune system, eventually becoming fully resistant. From a game theory perspective, this is the process of the virus “learning to play the game”. To simulate the space of possible viral escape trajectories against a chosen antibody, we employ an evolutionary algorithm. This approach approximates the process of viral evolution, where mutations that increase fitness (in our case, the virus’s payoff $R_{v}$) are more likely to persist, propagate, and become an ancestor of future generations.
We simulate this process of viral escape for a specified horizon of $H$ steps, with each step representing a single generation of viral evolution. This is the target duration of therapy efficacy. For each independent viral trajectory, we use the initial virus strain $v$ to generate a population of $P$ virus variants through random mutations. These are point mutations in the amino acid sequence of the viral antigen and they are introduced according to a predefined mutation rate. Each variant’s fitness is evaluated based on its payoff $R_{v}$ against the fixed antibody $a$. Importantly, this process captures the myopic nature of viral evolution: the virus optimises its fitness only with respect to the current antibody, unable to anticipate future human interventions or antibody changes. The next generation of the viruses is then created by probabilistically selecting the variants based on their fitness, with fitter variants more likely to be chosen for replication and subsequent mutation. This process repeats for $H$ generations, creating one trajectory of viral escape, see the Simulated Viral Escape via Evolution in Figure 1a and Methods section 4.2 for more details.
We use the antigen protein from the Dengue Virus as the virus for all our experiments, specifically, the structure with Protein Data Bank (PDB) code 2R29 [21, 22]. We discretise this structure using the Absolut! framework to generate the docking poses, which we then use to calculate the binding strength $B$ of any antibody-virus pair using our JAX implementation. In the viral escape step, we mutate only the amino acid sequence of the dengue envelope antigen, which is composed of 97 amino acids, and do not consider other components of Dengue Virus. Importantly, while we alter the amino acid sequence, we assume the structure of the antigen does not significantly change for the calculation of binding strength.
@@ 2.1.3 Optimisation of Antibody Shapers
Our antibody optimisation process aims to generate antibodies that can prevent viral escape. In particular, we desire antibodies which maintain their performance even once the virus has had the opportunity to evolve in response to them. For this purpose, we define the antibody fitness function $F^{H}_{v}(a)$ as the antibody’s performance over multiple potential viral escape trajectories of virus $v$:
@@
We then employ an evolutionary algorithm on the antibodies themselves to optimise this fitness function. In particular, optimisation proceeds iteratively for $N$ steps where we select the antibody with the highest fitness at each step and generate a new population of $P_{a}$ antibodies by introducing single mutations to the selected antibody. This process of evaluation, selection, and mutation continues, gradually improving the antibody’s robustness against, and ability to limit, viral escape, see Antibody Optimisation Loop in Figure 1a and Methods section 4.3 for more details.
We refer to the resulting optimised antibodies as shapers, or more specifically as $H=5$ shapers or $H=100$ shapers, reflecting the horizon length chosen for the viral escape simulations during their optimisation. Myopic antibodies are antibodies where horizon length $H=0$ and thus $F^{0}_{v}(a)=R_{a}(v,a)$. As we will discuss in more detail, the choice of $H$ significantly impacts the antibody’s long-term effectiveness at the cost of computational resources required for optimisation [^1] A single step of antibody optimisation requires simulating $P_{a}\times\eta$ independent viral escapes, each for horizon $H$ steps. However, we can parallelise over both $P_{a}$ and $\eta$., making horizon $H$ a vital parameter to consider.
From a machine learning perspective, this process can be viewed as meta-optimisation or meta-learning. The inner loop of our optimisation is the viral escape simulation, where we model the virus learning to evade the current antibody. The outer loop or meta-loop is the antibody optimisation process itself, which learns to generate antibodies that perform well across many possible viral escape trajectories. We are designing antibodies that anticipate and influence the escape process of the virus, embodying the core principle of opponent shaping in this biological context.
@@ 2.2 Antibody Shapers Minimise Viral Escape
We validate the effectiveness of the antibody shapers in optimising the escape-averaged antibody fitness function $F^{H}_{v}(a)$ compared to myopic antibodies that only respond to the current virus $v$. For our shaper antibodies, we select a long horizon of $H=100$ to capture extended viral escape trajectories. Both shapers and myopic antibodies are optimised for $N=30$ steps. Figure 2a presents the performance distributions of shapers and myopic antibodies under both objective functions.
@@ 2.2.2 Antibody Shapers with Varying Horizons
However, the number of antibody optimisation steps does not accurately reflect the computational or experimental cost of optimisation. Each simulation of viral escape requires a number of binding samples that increases linearly with the horizon length $H$. Yet, shorter horizon antibodies optimise an objective that diverges further from our “true” antibody objective $F^{100}_{v}(a)$. Due to this trade-off, we observe that the optimal training horizon varies depends on the available computational budget.
To illustrate this trade-off, we conduct an additional experiment shown in Figure 2d. Here, instead of fixing the number of optimisation steps $N$, we constrain the total number of binding samples — queries to our binding strength simulator used to evaluate all antibody and virus payoffs throughout the optimisation process — to be constant across different horizons. This approach provides a performance comparison that accounts for the computational resources necessary across varying horizon lengths. Interestingly, $H=20$ shapers perform strongly - nearly matching the performance of those optimised with horizon $H=100$ for a given number of antibody optimisation steps, and far exceeding it when accounting for the differing computational cost. This suggests that using a cheaper, shorter-horizon, proxy for the true antibody objective $F^{100}_{v}(a)$ can yield substantial benefits.
More generally, we find that the optimisation horizon significantly influences the performance of antibody shapers. While longer horizons generally lead to better long-term performance, the optimal horizon length is dependent on the available computational resources. Thus, it is important to consider the balance between computational cost and the fidelity of the optimisation objective when designing antibodies for long-term effectiveness against evolving viral populations.